# Score-based Generative Diffusion Models for Social Recommendations

### v1

@@ III-E Model Investigation
(1) Memory Analysis. We conducted additional experiments to assess the memory consumption of the social denoising models on the Ciao and Epinions dataset, employing consistent settings (batch size: 1024, embedding size: 128, 3-layer GNN encoder for both social embeddings and collaborative embeddings). The Table  V, reported in MiB, demonstrates that while memory efficiency is not a particular strength of SGSR among social denoising methods, memory usage remains within acceptable bounds.
TABLE V: Memory Analysis
@@
DSL 2942 13949
GBSR 3814 24330
SHaRe 3101 11993
SGSR 3299 15643
TABLE VI: Time Analysis
@@
(2) Time Analysis. We record the training time for various social denoising methods on three datasets, using consistent hyperparameter settings (batch size: 1024, embedding size: 128, learning rate: 0.0001, optimizer: Adam, 3-layer GNN encoder for both social embeddings and collaborative embeddings). The results are presented in table  VI. Despite leveraging the 5-step diffusion process, SGSR’s time consumption remains high due to the stepwise denoising. Besides, the training of SGSR learns the distribution of social representations rather than merely estimating user similarity. Future work may focus on achieving one-step embedding generation with a consistent model to reduce time costs.
(3) Scalability Analysis. In SGSR, the SGM is applied to the social representations of users, initially learned by GNN-based encoders. Unlike diffusion methods on graphs, the computational cost of the SGM does not increase with the graph size  [55]. Therefore, the scalability of SGSR is not constrained by graph size, particularly due to the diffusion model. This also motivates the idea of addressing the low-homophily challenge from a generative perspective rather than graph editing methods.
@@ IV Related Work
In this section, we summarize the related works on the social recommendation task and diffusion models for recommendations.
@@ IV-A Social Recommendations and Social Denoising Methods
(1) General Social Recommendation Methods. The widespread use of online social networks has alleviated the data sparsity issue in recommender systems by incorporating social context to enhance user representation modelling. Early approaches to social recommendations adopted the social homophily assumption and directly leveraged raw social networks, evolving through matrix factorization [56, 49, 57], graph neural networks [32, 12, 58], and self-supervised learning [50, 46, 17]. However, these methods struggle to address the redundant or irrelevant information inherent in raw social networks, introducing noise into user representations. This problem necessitates systematic denoising methods to improve social recommendation.
@@
(2) Data-centric Social Denoising Methods. Current social denoising methods primarily focus on modifying users’ social relations by assessing the confidence of social connections using handcrafted metrics to filter out irrelevant relations. For example, GDMSR [18] employs a transformer layer to evaluate the confidence of social connections based on historical interactions, removing a fixed percentage of edges with the lowest similarity scores. SHaRe [11] and MADM [59] use cosine similarity to assess the reliability of social relations, discarding or rewiring connections with low similarity scores and keeping reliable connections. GBSR [20] introduces a theoretically-inspired framework to maximize the mutual information between the denoised social graph and the interaction matrix, but it only drops the links in the social network while ignoring possibly beneficial edge rewiring or injecting operations. Generally speaking, existing data-centric social denoising methods can hardly reach the optimal social representation, since the best-fitted measure of social relations remains unknown and cannot be simply designed by human heuristics. Besides, graph editing methods only improve users’ social representations indirectly, and the editing path to the optimal graph structure corresponding to the optimal social representation may be computationally infeasible. This shows that it is necessary to address the low social homophily issue from an innovative generative perspective, which is the key research aim of this paper.
@@
(3) Contrastive Social Denoising Methods. Unlike data-centric methods that modify the structure of social networks to enhance user representations, another line of research focuses on directly improving users’ representations through contrastive learning. These methods [53, 60, 59] typically align user representations from different views (e.g., social or collaborative domains) to achieve a position more consistent with collaborative signals in the latent space. However, contrastive denoising approaches also rely on pre-defined user similarity measurements, which may not fully leverage the supervision signal from collaborative filtering, potentially falling short of reaching the optimal user representations. Additionally, since contrastive denoising methods adjust the training loss function to reduce noise in social networks, they are orthogonal to our proposed framework and can be adaptively integrated. We plan to explore the integration of the proposed SGSR framework with various auxiliary loss functions in future work.
@@ IV-B Diffusion Models for Recommendations
Diffusion models, a powerful class of generative models, have significantly reshaped various recommendation tasks with their effective generation capabilities. In collaborative filtering (CF), DiffRec [51] serves as a fundamental contribution, providing accurate recommendations from a diffusion perspective. Specifically, DiffRec encodes a user’s list of interacted items into a latent vector and employs a forward-reverse process to generate missing user-item interactions in the latent space. The diffused latent vector is then decoded into predicted user-item interactions. Following the proposal of the DiffRec framework, CF-Diff [61] and GiffCF [62] enhanced this approach by incorporating high-order relational capabilities through graph-level diffusion models. Another notable diffusion model, DreamRec [52], targets sequential recommendation tasks by using a conditional generation paradigm to generate target-item embeddings based on historical interaction sequences. Additionally, leveraging the strong generative ability of diffusion models to produce high-quality samples or representations for recommendation has shown promise in recommendations [63, 64, 55]. For instance, DDRM [55] refines noisy implicit user-item interaction data by applying a diffusion process to user and item representations, conditioned on their initial embeddings. Moreover, other tasks, such as point-of-interest (POI) recommendation [65] and knowledge-aware recommendation [66], have also benefited from the success of diffusion models. However, existing diffusion recommender models are not designed to integrate denoised social information into recommendations, which differs from the primary research objective of this paper. Additionally, most existing works utilize the Denoising Diffusion Probabilistic Models (DDPM) [28] framework, which can be seen as a specific discretized form of Score-based Generative Models (SGM) [30], highlighting the novelty of the proposed SGSR framework.
@@ V Conclusion
For future work, a promising direction is to accelerate sampling by designing consistent functions [67] within the recommendation context (training time analysis in Appendix  VI). Additionally, exploring various SDE solvers, such as Runge-Kutta methods, could enhance the inference process by leveraging the computational flexibility of score-based diffusion models. Furthermore, while this paper primarily focuses on existing VP-SDEs and VE-SDEs, it is worthwhile to investigate more expressive and suitable SDEs tailored to the recommendation context.

### v2

@@ III-E Model Investigation
(1) Memory Analysis. We conducted additional experiments to assess the memory consumption of the social denoising models on the Ciao and Epinions datasets, employing consistent settings (batch size: 1024, embedding size: 128, 3-layer GNN encoder for both social embeddings and collaborative embeddings). The Table  VI, reported in MiB, demonstrates that while memory efficiency is not a particular strength of SGSR among social denoising methods, memory usage remains within acceptable bounds.
TABLE VI: Memory Analysis
@@
DSL 2,942 13,949
GBSR 3,814 24,330
SHaRe 3,101 11,993
SGSR 3,299 15,643
TABLE VII: Time Analysis
@@
(2) Time Analysis. We record the training time for various social denoising methods on three datasets, using consistent hyperparameter settings (batch size: 1024, embedding size: 128, learning rate: 0.0001, optimizer: Adam, 3-layer GNN encoder for both social embeddings and collaborative embeddings). The results are presented in Table VII. Despite leveraging the 5-step diffusion process, SGSR’s time consumption remains high due to the stepwise denoising. Besides, the training of SGSR learns the distribution of social representations rather than merely estimating user similarity. Future work may focus on achieving one-step embedding generation with a consistent model to reduce time costs.
(3) Scalability Analysis. In SGSR, the SGM is applied to the social representations of users, initially learned by GNN-based encoders. Unlike diffusion methods on graphs, the computational cost of the SGM does not increase with the graph size  [50]. Therefore, the scalability of SGSR is not constrained by graph size, particularly due to the continuous-state diffusion model. This also motivates the idea of addressing the low-homophily challenge from a generative perspective rather than graph editing methods.
@@
III-F Statistical Analyses
To further validate the effectiveness of SGSR, we statistically analyze the learned user representations from the Ciao dataset to evaluate their capacity for capturing real social connections. Specifically, we compute pairwise cosine similarities between user embeddings and select the top-$K$ most similar users as predicted social connections. These predictions are then quantitatively assessed against first-order ground-truth relationships using Recall@$K$ (R@$K$) and Hit Ratio@$K$ (HR@$K$) metrics, where $K$ is in {20, 10, 5}. For comparative analysis, we examine representations from GDSSL, which combines graph diffusion with self-supervised learning for social denoising, alongside LightGCN as a pure collaborative filtering baseline. Table VIII demonstrates that both social denoising models significantly outperform LightGCN, with SGSR achieving consistent improvements over GDSSL. These results confirm that: (1) social denoising models effectively learn social connection patterns, and (2) SGSR’s denoised representations better preserve social relationship structures than GDSSL’s approach.
TABLE VIII: Statistical Analyses of Social Connections
Method R@20 R@10 R@5 HR@20 HR@10 HR@5
LightGCN 0.0213 0.0133 0.0086 0.2667 0.1805 0.1996
GDSSL 0.1866 0.1425 0.1030 0.7996 0.7056 0.5839
SGSR 0.1996 0.1513 0.1122 0.8234 0.7277 0.6269
@@ IV Related Work
This section reviews related work in social recommendation and diffusion models for recommender systems.
@@ IV-A Social Recommendations and Social Denoising Methods
(1) General Social Recommendation Methods. The widespread use of online social networks has alleviated the data sparsity issue in recommender systems by incorporating social context to enhance user representation modelling. Early approaches to social recommendations adopted the social homophily assumption and directly leveraged raw social networks, evolving through matrix factorization [40, 51], graph neural networks [28, 11, 52], and self-supervised learning [41, 37, 16]. However, these methods struggle to address the redundant or irrelevant information inherent in raw social networks, introducing noise into user representations. This problem necessitates systematic denoising methods to improve social recommendation.
@@
(2) Data-centric Social Denoising Methods. Current social denoising methods primarily focus on modifying users’ social relations by assessing the confidence of social connections using handcrafted metrics to filter out irrelevant relations. For example, GDMSR [17] employs a transformer layer to evaluate the confidence of social connections based on historical interactions, removing a fixed percentage of edges with the lowest similarity scores. SHaRe [10] and MADM [53] use cosine similarity to assess the reliability of social relations, discarding or rewiring connections with low similarity scores and keeping reliable connections. GBSR [19] introduces a theoretically-inspired framework to maximize the mutual information between the denoised social graph and the interaction matrix, but it only drops the links in the social network while ignoring possibly beneficial edge rewiring or injecting operations. Generally speaking, existing data-centric social denoising methods can hardly reach the optimal social representation, since the best-fitted measure of social relations remains unknown and cannot be simply designed by human heuristics. Besides, graph editing methods only improve users’ social representations indirectly, and the editing path to the optimal graph structure corresponding to the optimal social representation may be computationally infeasible. This highlights the necessity for innovative generative solutions to low social homophily, which is the central focus of our study.
@@
(3) Contrastive Social Denoising Methods. Unlike data-centric methods that modify the structure of social networks to enhance user representations, another line of research focuses on directly improving users’ representations through contrastive learning. These methods [46, 53] typically align user representations from different views (e.g., social or collaborative domains) to achieve a position more consistent with collaborative signals in the latent space. However, contrastive denoising approaches also rely on pre-defined user similarity measurements, which may not fully leverage the supervision signal from collaborative filtering, potentially falling short of reaching the optimal user representations. Additionally, since contrastive denoising methods adjust the training loss function to reduce noise in social networks, they are orthogonal to our proposed framework and can be adaptively integrated.
@@ IV-B Diffusion Models for Recommendations
Diffusion models, a powerful class of generative models, have significantly reshaped various recommendation tasks with their effective generation capabilities. In collaborative filtering (CF), DiffRec [42] serves as a fundamental contribution, providing accurate recommendations from a diffusion perspective. Specifically, DiffRec encodes a user’s list of interacted items into a latent vector and employs a forward-reverse process to generate missing user-item interactions in the latent space. The diffused latent vector is then decoded into predicted user-item interactions. Following the proposal of the DiffRec framework, CF-Diff [54] and GiffCF [45] enhanced this approach by incorporating high-order relational capabilities through graph-level diffusion models. Another notable diffusion model, DreamRec [43], targets sequential recommendation tasks by using a conditional generation paradigm to generate target-item embeddings based on historical interaction sequences. Additionally, leveraging the strong generative ability of diffusion models to produce high-quality samples or representations for recommendation has shown promise in recommendations [55, 56, 50]. For instance, DDRM [50] refines noisy implicit user-item interaction data by applying a diffusion process to user and item representations, conditioned on their initial embeddings. However, existing diffusion recommender models are not designed to integrate denoised social information into recommendations, which differs from the primary research objective of this paper. Additionally, most existing works utilize the Denoising Diffusion Probabilistic Models (DDPM) [24] framework, which can be seen as a specific discretized form of Score-based Generative Models (SGM) [26], highlighting the novelty of the proposed SGSR framework.
@@ V Conclusion
For future work, a promising direction is to accelerate sampling by designing consistent functions [57] tailored to the recommendation context to enable one-step inference. Additionally, exploring various SDE solvers, such as Runge-Kutta methods, could further enhance the inference efficiency by leveraging the computational flexibility of score-based diffusion models. From a training perspective, investigating model distillation techniques to efficiently train a student model guided by SGSR along the distribution trajectory also presents a valuable direction.
Acknowledgments
The research described in this paper has been partially supported by the National Natural Science Foundation of China (project no. 62102335), General Research Funds from the Hong Kong Research Grants Council (project no. PolyU 15207322, 15200023, 15206024, and 15224524), internal research funds from Hong Kong Polytechnic University (project no. P0042693, P0048625, and P0051361). This work was supported by computational resources provided by The Centre for Large AI Models (CLAIM) of The Hong Kong Polytechnic University.